given that cos(2sin inverse(x))=1/9
we have find  the value of x=?

Grade:12

1 Answers

Aditya Gupta
5 years ago
let sin inverse(x)= y so that x= siny
given cos2y= 1/9
or 1 – 2sin^2y = 1/9
or 1 – 2x^2 = 1/9
or x^2= 4/9
or x= +-(2/3)